# Private Counterfactual Retrieval

### v1

@@ Abstract
Transparency and explainability are two extremely important aspects to be considered when employing black-box machine learning models in high-stake applications. Providing counterfactual explanations is one way of catering this requirement. However, this also poses a threat to the privacy of both the institution that is providing the explanation as well as the user who is requesting it. In this work, we propose multiple schemes inspired by private information retrieval (PIR) techniques which ensure the user’s privacy when retrieving counterfactual explanations. We present a scheme which retrieves the exact nearest neighbor counterfactual explanation from a database of accepted points while achieving perfect (information-theoretic) privacy for the user. While the scheme achieves perfect privacy for the user, some leakage on the database is inevitable which we quantify using a mutual information based metric. Furthermore, we propose strategies to reduce this leakage to achieve an advanced degree of database privacy. We extend these schemes to incorporate user’s preference on transforming their attributes, so that a more actionable explanation can be received. Since our schemes rely on finite field arithmetic, we empirically validate our schemes on real datasets to understand the trade-off between the accuracy and the finite field sizes.
1 INTRODUCTION
@@
With the growing call for the right to an explanation (Voigt and Bussche, 2017; Park, 2023), the framework of counterfactual explanations has generated immense interest as a means to explain the decision-making of complex models in high-stakes applications (Wachter et al., 2017). Counterfactual explanations provide the minimum input perturbation required to alter a model outcome, and are closely tethered to algorithmic recourse, e.g., increase your income by 10K to qualify for a loan.
@@
Counterfactual explanations are also susceptible to privacy concerns. For instance, (Pawelczyk et al., 2023; Yang et al., 2022) bring out privacy issues related to the underlying training data, while (Aïvodji et al., 2020; Wang et al., 2022; Dissanayake and Dutta, 2024) study model extraction using counterfactual explanations. However, these works predominantly focus on privacy from the institution’s side. Applicant privacy concern arises if they wish to obtain their counterfactual explanations privately without revealing their current input feature vector to the institution.
@@
An applicant may be reluctant to share their entire feature vector with the institution for several reasons, e.g., a formal application process might be expensive in terms of time and resources, or allow for only a limited number of attempts (Dissanayake and Dutta, 2024), or they might wish to preserve the privacy of their data until they improve their chances of acceptance.
@@
Our work introduces the novel problem of private counterfactual retrieval (PCR). The objective of PCR is to design a strategy that applicants and institutions can jointly agree upon that enables: (i) the applicant to privately retrieve their counterfactual explanation from an institution through an alternate set of queries; and (ii) the institution to also not leak any further information beyond what the applicant requires. To this end, we draw inspiration from the problem of private information retrieval (PIR) (Chor et al., 1998; Ulukus et al., 2022) which enables a user to download a message from a set of messages stored in a system of databases without revealing the index of the desired message. Along the lines of PIR, we assume that the institution has a stored database of accepted applicants, and the database entries lie in a finite field. We seek to retrieve the index of the exact nearest neighbor for an applicant from the database without revealing their own feature vector (in an information-theoretic sense).
@@
    While our strategies achieve perfect privacy, we also perform an empirical analysis to understand the implications of our finite field assumptions on real data. We compare the accuracy loss from translating real-valued data to finite-field data to assure that the designed schemes act as intended, and to understand the trade-off between the field size requirements and the efficacy of our schemes.
1.1 RELATED WORKS
@@
Counterfactual explanations and privacy: Since the initial formulation in Wachter et al. (2017), numerous works have been focusing on generating counterfactual explanations with different properties. Proximity to the query instance (Brughmans et al., 2023), robustness (Upadhyay et al., 2021; Hamman et al., 2023), actionability (Poyiadzi et al., 2020; Mothilal et al., 2019), sparsity in change (Brughmans et al., 2023; Mothilal et al., 2019), and diversity (Mothilal et al., 2019) are some such properties; we refer the reader to Karimi et al. (2022) and Guidotti (2024) for a comprehensive survey on different methods. In this work, we focus on both proximity to the original instance and actionability. Moreover, we use nearest-neighbor counterfactuals as the explanation method considered. Existing works on privacy within the context of counterfactual explanations mainly focus on the institution’s end. In this regard, Pawelczyk et al. (2023) analyzes inferring the membership of an explanation in the training set of the model while Yang et al. (2022) provides differentially-private counterfactuals. Aïvodji et al. (2020) and Wang et al. (2022) present two ways of utilizing counterfactuals to extract the model when counterfactuals are provided for any query. Dissanayake and Dutta (2024) presents a model extraction strategy that specifically leverages the fact that the counterfactuals are close to the decision boundary. Explanation linkage attacks that try to extract private attributes of a nearest-neighbor counterfactual explanation are discussed in Goethals et al. (2023). All of these works focus on the privacy of either the model or the data stored in the institution’s database. In contrast, we are interested in the privacy of the applicant who is asking for an explanation.
@@
Private information retrieval: The PIR problem formulation was first defined in Chor et al. (1998) and its capacity, i.e., the maximum ratio between the number of required message symbols and the number of total downloaded symbols, was found in Sun and Jafar (2017). Different from the original formulation which required non-colluding databases with replicated contents, other variants, such as PIR with colluding databases (Sun and Jafar, 2018a; Yao et al., 2021), PIR with coded databases, (Tajeddine et al., 2018; Banawan and Ulukus, 2018) were also considered. In SPIR, an extra requirement is that the user cannot get any information beyond its required message. The capacity of SPIR was found in Sun and Jafar (2018b) and the variants with colluding and coded databases appeared in Wang and Skoglund (2019). Reference Jia et al. (2019), proposes a cross subspace alignment (CSA) approach as a unifying framework for PIR and SPIR with additional requirements, such as security against the storing units, e.g., servers. These schemes are capacity-achieving in some cases, for instance, when the number of messages is large. We refer the reader to Ulukus et al. (2022) for a comprehensive survey on the PIR and SPIR literature.
@@
Private approximate nearest neighbor search: Closely related to our problem is the nearest neighbor search problem, where the user needs to retrieve the indices of the vectors in a database, that are closest to their vector according to some similarity metric. In this regard, Servan-Schreiber et al. (2022) proposed algorithms that guaranteed computational privacy, both to the user and the database, while the user retrieves a sub-optimal nearest neighbor. Reference Vithana et al. (2024) proposed an information-theoretically private clustering-based solution based on the dot-product metric. This work, however, did not consider database privacy, and the user retrieves only an approximate nearest neighbor.
2 SYSTEM MODEL
@@
We assume a pre-trained two-class classification model that takes as input a $d$-dimensional feature vector and classifies it into its target class, e.g., accepted or rejected. A user who is rejected by this model wishes to privately retrieve a valid counterfactual sample corresponding to their data sample. However, the user does not have access to the model and relies on a database $\mathcal{D}$ that contains the feature vectors of a set of samples accepted by the same model as depicted in Figure 1. We assume that each attribute of the samples is an integer in $[0:R]$. The samples in $\mathcal{D}$ are stored at a server and are indexed as $y_{1},y_{2},\ldots,y_{M}$ where $M=|\mathcal{D}|$. The samples are stored in $N$ non-colluding and non-communicating servers in a replicated manner. The goal of the user is to retrieve the index of the accepted data sample that is nearest to their sample $x$ based on a preference vector [^1] We consider the system model based on actionable vector $w$ as the general system model. $w$ which is globally known. Our system can be simplified to the case of equal actionability on all features by choosing $w=\mathbf{1}^{t}$., $w$, according to some distance metric $d_{w}(.,.)$, i.e.,
@@
To quantify privacy for the servers, we consider an information-theoretic measure that defines the amount of information leakage about the samples in $\mathcal{D}$ to the user upon receiving the answers as follows,
@@
The requirements stated above can be formulated as an optimization problem as follows
$\min_{\mathcal{A},\mathcal{Q},\mathcal{S}}\quad$ $I(y_{1},\ldots,y_{M};Q_{[N]}^{[x,w]},A_{[N]}^{[x,w]}|x,w)$ (6)
s.t. $H(\theta^{*}|A_{[N]}^{[x,w]},x,w)=0$ (7)
$I(x,w,\theta^{*};Q_{n}^{[x,w]},A_{n}^{[x,w]}|y_{1},\ldots,y_{M})=0$ (8)
where $\mathcal{A}$ is the set of all possible answers the servers can send, $\mathcal{Q}$ are all the possible queries the user can send, and $\mathcal{S}$ are all the possible ways to encode and store the accepted samples in the $N$ servers.
In terms of distance metric, we use the (weighted) $\ell_{2}$ norm, i.e., $d_{w}(u,v)=(u-v)^{T}\bm{W}(u-v)$ where $\bm{W}$ is the diagonal matrix with $w$ on its main diagonal. Our results can be extended easily using any $\ell_{n}$ norm with $n$ being even, as well as to the dot product metric [^2] It is important to note that if dot product metric is used, actionability can be implemented directly in the rejected sample sent by the user.. For space limitations, we define Vandermonde matrices on the finite field $\mathbb{F}_{q}$ where $q$ is a positive prime power.
@@
$\bm{M}_{n}(i,j)=\begin{cases}1,&j=1,\\ \alpha_{i}^{j-1},&j=2,\ldots,n,\quad i=1,\ldots,n.\end{cases}$ (9)
@@
In addition, for completeness one-time pad theorem, which we use extensively, is stated below as in Cover and Thomas (1999).
Theorem 1 (One-time pad theorem)
Let $q$ be a prime, then for any random variable $X\in\mathbb{F}_{q}$, and any uniform random variable $Z$ chosen at random in $\mathbb{F}_{q}$
$I(X;X+Z)=0,$ (10)
where addition is according to $\mathbb{F}_{q}$ arithmetic.
We divide the description of our schemes into two sections, Section 3 without user preference and Section 4 with private user actionability.
3 PROPOSED PCR SCHEMES
In this section, we assume $w=\mathbf{1}^{t}$, thus $d_{w}(.,.)$ is the squared $\ell_{2}$ norm. Consequently, we simplify the notations of queries and answers for server $n$ to $Q_{n}^{[x]}$ and $A_{n}^{[x]}$, respectively. For ease of exposition, we write
@@ 3.1 Baseline PCR
There exists a scheme that can retrieve the exact closest counterfactual index from $N=2$ servers with perfect user privacy. In addition, the communication cost for the scheme is $2(d+M)$ symbols.
To prove the achievability of Theorem 2, we present our baseline PCR scheme. Let the operating field be $\mathbb{F}_{q}$ where $q$ is a prime satisfying $q>R^{2}d$. Each server stores the $M$ accepted samples, $y_{1},\ldots,y_{M}$, each being a $d$ dimensional vector. Let $\alpha_{1},\alpha_{2}$ be two distinct elements of $\mathbb{F}_{q}$ globally known to the user and servers. The user privately generates a random vector $Z$ uniformly from $\mathbb{F}_{q}^{d}$ and then sends the query $Q_{n}^{[x]}$ to the $n$th server based on $x$ as follows
@@
where $\alpha_{1},\alpha_{2}$ are distinct elements in the field and globally known constants. Since $x$ is one-time padded with $\alpha_{n}Z$ which is uniform on $\mathbb{F}_{q}$, each server learns no information on $x$, i.e., $I(x;Q_{n}^{[x]}|y_{1},\ldots,y_{M})=0$. Let the servers share $M$ independent random variables $Z^{\prime}(i),i\in[M]$ picked uniformly from $\mathbb{F}_{q}$. Given a query, the servers compute one answer for each $y_{i}\in\mathcal{D}$ as follows
@@
Note that $A_{n}^{[x]}(i)$ is a second degree polynomial of $\alpha_{n}$ with $\alpha_{n}^{2}||Z||^{2}$ already known to the user. Therefore, the user cancels $\alpha_{1}^{2}||Z||^{2}$ and $\alpha_{2}^{2}||Z||^{2}$ from $A_{1}^{[x]}(i)$ and $A_{2}^{[x]}(i)$, respectively, to obtain
@@
3.2 Diff-PCR
@@
There exists a solution for the optimization problem defined in (6) with a lower value for the objective function, i.e., better privacy to the servers, than the baseline PCR. In addition, the communication cost of the scheme is $2(d+M-1)$.
@@
The servers share a common random vector $Z^{\prime}_{[M-1]}=[Z^{\prime}(1),Z^{\prime}(2),\ldots,Z^{\prime}(M-1)]^{t}$, where each entry is picked uniformly and independently from $\mathbb{F}_{q}$. As described in Section 3.1, the user sends the query given in (12) to server $n=1,2$. Then, server $n$ computes the following answer for each $i\in[M-1]$,
@@
Finding $y_{i}$ Closest to $x$:
The user finds the index $\theta^{*}$ of their counterfactual using Algorithm 1.
@@
Leakage Analysis:

### v2

@@ Abstract
Transparency and explainability are two extremely important aspects to be considered when employing black-box machine learning models in high-stake applications. Providing counterfactual explanations is one way of fulfilling this requirement. However, this also poses a threat to the privacy of both the institution that is providing the explanation as well as the user who is requesting it. In this work, we propose multiple schemes inspired by private information retrieval (PIR) techniques which ensure the user’s privacy when retrieving counterfactual explanations. We present a scheme which retrieves the exact nearest neighbor counterfactual explanation from a database of accepted points while achieving perfect (information-theoretic) privacy for the user. While the scheme achieves perfect privacy for the user, some leakage on the database is inevitable which we quantify using a mutual information based metric. Furthermore, we propose strategies to reduce this leakage to achieve an advanced degree of database privacy. We extend these schemes to incorporate user’s preference on transforming their attributes, so that a more actionable explanation can be received. Since our schemes rely on finite field arithmetic, we empirically validate our schemes on real datasets to understand the trade-off between the accuracy and the finite field sizes. Finally, we present numerical results to support our theoretical findings, and compare the database leakage of the proposed schemes.
I INTRODUCTION
@@
With the growing call for the right to an explanation [1, 2], the framework of counterfactual explanations has generated immense interest as a means to explain the decision-making of complex models in high-stakes applications [3]. Counterfactual explanations provide the minimum input perturbation required to alter a model outcome, and are closely tethered to algorithmic recourse, e.g., increase your income by 10K to qualify for a loan.
@@
Counterfactual explanations are also susceptible to privacy concerns. For instance, [4, 5] bring out privacy issues related to the underlying training data, while [6, 7, 8] study model extraction using counterfactual explanations. However, these works predominantly focus on privacy from the institution’s side. Applicant privacy concern arises if they wish to obtain their counterfactual explanations privately without revealing their current input feature vector to the institution.
@@
An applicant may be reluctant to share their entire feature vector with the institution for several reasons, e.g., a formal application process might be expensive in terms of time and resources, or allow for only a limited number of attempts [8], or they might wish to preserve the privacy of their data until they improve their chances of acceptance.
@@
Our work introduces the novel problem of private counterfactual retrieval (PCR). The objective of PCR is to design a strategy that applicants and institutions can jointly agree upon that enables: (i) the applicant to privately retrieve their counterfactual explanation from an institution through an alternate set of queries; and (ii) the institution to also not leak any further information beyond what the applicant requires. To this end, we draw inspiration from the problem of private information retrieval (PIR) [9, 10] which enables a user to download a message from a set of messages stored in a system of databases without revealing the index of the desired message. Along the lines of PIR, we assume that the institution has a stored database of accepted applicants, and the database entries lie in a finite field. We seek to retrieve the index of the exact nearest neighbor for an applicant from the database without revealing their own feature vector (in an information-theoretic sense). This yields a new and complementary formulation and renders our work incomparable to the prior works since they do not attain perfect information-theoretic privacy.
@@
    We also perform an empirical analysis to understand the implications of our finite field assumptions on real data. We compare the accuracy loss from translating real-valued data to finite-field data to ensure that the designed schemes act as intended, and to understand the trade-off between the field size requirements and the efficacy of our schemes.
  5. 5.
    Finally, we numerically evaluate the database leakage for the proposed PCR schemes, and verify the comparative leakage values among the proposed schemes.
I-A RELATED WORKS
Counterfactual explanations and privacy
@@ I-A RELATED WORKS
Since the initial formulation in [3], numerous works have been focusing on generating counterfactual explanations with different properties. Proximity to the query instance [11], robustness [12, 13], actionability [14, 15], sparsity in change [11, 15], and diversity [15] are some such properties; we refer the reader to [16] and [17] for a comprehensive survey on different methods. In this work, we focus on both proximity to the original instance and actionability. Moreover, we use nearest-neighbor counterfactuals as the explanation method considered. Existing works on privacy within the context of counterfactual explanations mainly focus on the institution’s end. In this regard, [4] analyzes inferring the membership of an explanation in the training set of the model while [5] provides differentially-private counterfactuals. [6] and [7] present two ways of utilizing counterfactuals to extract the model when counterfactuals are provided for any query. [8] presents a model extraction strategy that specifically leverages the fact that the counterfactuals are close to the decision boundary. Explanation linkage attacks that try to extract private attributes of a nearest-neighbor counterfactual explanation are discussed in [18]. All of these works focus on the privacy of either the model or the data stored in the institution’s database. In contrast, we are interested in the privacy of the applicant who is asking for an explanation.
Private information retrieval
@@
The PIR problem formulation was first defined in [9] and its capacity, i.e., the maximum ratio between the number of required message symbols and the number of total downloaded symbols, was found in [19]. Different from the original formulation which required non-colluding databases with replicated contents, other variants, such as PIR with colluding databases [20, 21], PIR with coded databases, [22, 23] were also considered. In SPIR, an extra requirement is that the user cannot get any information beyond its required message. The capacity of SPIR was found in [24] and the variants with colluding and coded databases appeared in [25]. Reference [26], proposes a cross-subspace alignment (CSA) approach as a unifying framework for PIR and SPIR with additional requirements, such as security against the storing units, e.g., servers. These schemes are capacity-achieving in some cases, for instance, when the number of messages is large. We refer the reader to [10] for a comprehensive survey on the PIR and SPIR literature.
@@
Private approximate nearest neighbor search: Closely related to our problem is the nearest neighbor search problem, where the user needs to retrieve the indices of the vectors in a database, that are closest to their vector according to some similarity metric. In this regard, [27] proposed algorithms that guarantee computational privacy, both to the user and the database, while the user retrieves a sub-optimal nearest neighbor. Reference [28] proposed an information-theoretically private clustering-based solution based on the dot-product metric. This work, however, did not consider database privacy, and the user retrieves only an approximate nearest neighbor.
II SYSTEM MODEL
@@
The institution has a pre-trained binary classification model that takes as input a $d$-dimensional feature vector and classifies it into its target class, e.g., accepted or rejected. A user who is rejected by this model wishes to privately retrieve a valid counterfactual sample corresponding to their data sample. However, the user does not have access to the model and relies on a database $\mathcal{D}$ that contains the feature vectors of a set of samples accepted by the model as depicted in Fig. 1. We assume that each attribute of the samples is an integer in $[0:R]$. The samples in $\mathcal{D}$ are stored in $N$ non-colluding and non-communicating servers in a replicated manner. The samples are indexed as $y_{1},y_{2},\ldots,y_{M}$ where $M=|\mathcal{D}|$. The goal of the user is to retrieve the index of the accepted data sample that is nearest to their sample $x$ based on a preference vector [^1] We consider the system model based on actionable vector $w$ as the general system model. $w$ which is globally known. Our system can be simplified to the case of equal actionability on all features by choosing $w=\mathbf{1}^{t}$., $w$, according to some distance metric $d_{w}(.,.)$, i.e.,
@@
To quantify privacy for the servers, we consider an information-theoretic measure that defines the amount of information leakage about the samples in $\mathcal{D}$ to the user, upon receiving the answers, as follows,
@@
Our goal is to develop schemes that reduce this database leakage, while maintaining perfect user privacy. We use the (weighted) $\ell_{2}$ norm, i.e., $d_{w}(u,v)=(u-v)^{T}\bm{W}(u-v)$ where $\bm{W}$ is the diagonal matrix with $w$ on its main diagonal. Our results can be extended easily using any $\ell_{n}$ norm with $n$ being even, as well as to the dot product metric [^2] It is important to note that if dot product metric is used, actionability can be implemented directly in the rejected sample sent by the user.. For space limitations, we define Vandermonde matrices on the finite field $\mathbb{F}_{q}$ where $q$ is a positive prime power.
@@
The benefit of privately retrieving the index of the closest neighbor is to guarantee that the institution does not learn any information about the user’s input. Further, it does not learn the counterfactual explanation corresponding to the user. This is achieved in two steps: First, we perform the PCR scheme to retrieve the counterfactual index from the server. Next, we use this index to perform the SPIR scheme [24] on the same dataset. The SPIR scheme applies seamlessly in the existing system due to the replicated and non-colluding databases.
@@
$\bm{M}_{n}(i,j)=\alpha_{i}^{j-1},\quad i,j\in[n].$ (6)
@@
We divide the description of our schemes into two sections, Section III without user preference and Section IV with private user actionability.
Finally, it is important to note that this formulation for counterfactual retrieval is novel, as far as we know. This is mainly because perfect user privacy is required and the servers accept some data leakage. Thus it is unfair to compare our model and schemes developed here with other counterfactual retrieval schemes.
III PROPOSED PCR SCHEMES
In this section, we assume $w=\mathbf{1}^{t}$, thus $d_{w}(.,.)$ is the squared $\ell_{2}$ norm. Consequently, we simplify the notations of queries and answers for server $n$ to $Q_{n}^{[x]}$ and $A_{n}^{[x]}$, respectively. For ease of exposition, we denote
@@ III-A Baseline PCR
There exists a scheme that can retrieve the exact closest counterfactual index from $N=2$ servers with perfect user privacy, with the communication cost of $2(d+M)$ symbols of $\mathbb{F}_{q}$ where $q>R^{2}d$ is prime.
To prove the achievability of Theorem 1, we present our baseline PCR scheme. Let the operating field be $\mathbb{F}_{q}$ where $q$ is a prime satisfying $q>R^{2}d$, where $d$ is the number of features for each accepted sample and $R$ is the maximum value each feature can have. In addition, we require $q$ to be greater than the aforementioned value to make sure that addition in this field will not reverse the order of the closeness of the samples with the user sample. Although fields can be defined using prime powers, we restrict ourselves to only primes since prime power fields, $\mathbb{F}_{q^{r}}$, are isomorphic to $\mathbb{F}_{q}^{r}$ which means that they are not ordered. Each server stores the $M$ accepted samples, $y_{1},\ldots,y_{M}$, each being a $d$ dimensional vector. Let $\alpha_{1},\alpha_{2}$ be two distinct elements of $\mathbb{F}_{q}$ globally known to the user and servers. The user privately generates a random vector $Z$ uniformly from $\mathbb{F}_{q}^{d}$ and then sends the query $Q_{n}^{[x]}$ to the $n$th server based on $x$ as follows,
@@
where $\alpha_{1},\alpha_{2}$ are distinct elements in the field and globally known constants. Since $x$ is one-time padded with $\alpha_{n}Z$ which is uniform on $\mathbb{F}_{q}$, each server learns no information on $x$, i.e., $I(x;Q_{n}^{[x]}|y_{1},\ldots,y_{M})=0$. Let the servers share $M$ independent random variables $Z^{\prime}(i),i\in[M]$ picked uniformly from $\mathbb{F}_{q}$. Given a query, the servers compute one answer for each $y_{i}\in\mathcal{D}$ as follows,
@@
Note that $A_{n}^{[x]}(i)$ is a second degree polynomial of $\alpha_{n}$ with $\alpha_{n}^{2}||Z||^{2}$ already known to the user. Therefore, the user cancels $\alpha_{1}^{2}||Z||^{2}$ and $\alpha_{2}^{2}||Z||^{2}$ from $A_{1}^{[x]}(i)$ and $A_{2}^{[x]}(i)$, respectively,
@@
Computation Complexity
At the user side, the computation for the queries involves two scalar multiplications and additions in $d$ dimensions, which is $O(d)$. To find $\theta^{*}$, the user decodes $d_{i}(x)$ using (10) for each $i\in[M]$ and compares them to find the minimum, both using $O(M)$ operations. Therefore, the user-side computational complexity is $O(d+M)$. Each server generates $M$ answers, where each answer involves subtraction, squaring, and addition of $d$ elements of each vector. This results in $O(Md)$ operations at each server.
Remark 1
It is important to note that (17) is evaluated prior to the scheme initiation and done once only till the dataset itself changes. This is a part of pre-processing for Mask-PCR and is not a part of the main scheme, which begins with the user sending their query. This is why it is not included in the calculation of the computation complexity.
III-B Diff-PCR
@@
There exists a scheme with a lower value of database leakage (5) compared to the baseline PCR, with the communication cost of $2(d+M-1)$ symbols of $\mathbb{F}_{q}$ where $q>2R^{2}d$ is prime.
@@
The servers share a common random vector $Z^{\prime}_{[M-1]}=[Z^{\prime}(1),Z^{\prime}(2),\ldots,Z^{\prime}(M-1)]^{t}$, where each entry is picked uniformly and independently from $\mathbb{F}_{q}$. As described in Section III-A, the user sends the query given in (8) to server $n=1,2$. Then, server $n$ computes the following answer for each $i\in[M-1]$,
@@
Finding $y_{i}$ Closest to $x$: The user finds the index $\theta^{*}$ of their counterfactual using Algorithm 1.
@@
Computation Complexity
The user side query generation is $O(d)$ as in the baseline scheme. To find $\theta^{*}$, the user finds $d_{i}(x)-d_{i+1}(x)$ using (14) and evaluates $\theta^{*}$ using Algorithm 1, which has a complexity of $O(M)$. Thus, the user’s computational complexity is $O(d+M)$. The server side complexity is $O(Md)$, same as the baseline.
Leakage Analysis